OpenStudy (lilai3):

Please help! Will award medal/fan/write testimony, etc. etc!! The Centerville HIgh School soccer team has a goal of winning at least 60% of their 18 games this season. In the first three weeks, the team has won 4 games. How many more games must the team win to meet their goal? Write the inequality and explain. Thank you!~

OpenStudy (agent0smith):

First you'll need to find what 60% of 18 games is.

OpenStudy (lilai3):

uhm...0.6(18)... wait why? and it's 10.8, by the way.

OpenStudy (agent0smith):

Because... "has a goal of winning at least 60% of their 18 games this season."

OpenStudy (agent0smith):

Think about it...

OpenStudy (lilai3):

oh..okay. never mind. now it seemed like a stupid question. sorry.

OpenStudy (agent0smith):

They've won 4. They need to win 10.8. How many more games must they win?

OpenStudy (lilai3):

so it's a subtraction problem here... 10.8-4=6.8 games... but how can anyone win 6 POINT 8 games?...

OpenStudy (agent0smith):

Well, then what should you do with the 6.8?

OpenStudy (lilai3):

...well, i don't know...round it?

OpenStudy (agent0smith):

Yes

OpenStudy (lilai3):

okay, 7?

OpenStudy (lilai3):

so what's the inequality?

OpenStudy (agent0smith):

Yes. And note that even if it was 6.1, you'd round UP to 7 - if they win 6 games, that's not enough, they must win 7 to have enough. Well, they need to win at least 7. So that means their number of wins from now on, must be greater than or equal to 7.

OpenStudy (agent0smith):

If x is the number of wins from now, then \[\large x \ge 7\]

OpenStudy (lilai3):

oh, i see...but what would be the inequality if one DIDN'T solve it? \[x+4 \ge 0.6\times18?\]

OpenStudy (agent0smith):

Yes

OpenStudy (lilai3):

okay thank you.

OpenStudy (lilai3):

The length of the base of the triangle is less than the height of the triangle. What are the possible values of x? Formulate a linear ineqaulity to solve the problem. Determine whether your answers are reasonable.

OpenStudy (lilai3):

inequality*** btw, please help me if you can

OpenStudy (agent0smith):

Base < Height

OpenStudy (lilai3):



OpenStudy (agent0smith):

Base < Height use this but put in the numbers on the diagram.

OpenStudy (lilai3):

12<4+x?

OpenStudy (agent0smith):

Yes

OpenStudy (lilai3):

so are the possible values to x: 9, 10, 11, 12,13 ,14, 15, 16, 17, 18, 19, 20, etc. etc. etc. all the way to INFINITAY?

OpenStudy (agent0smith):

lol yes, x>8 in other words

OpenStudy (lilai3):

ok thank you you have geen breat help... i meant you have BEEN GREAT help...ahem...excuse me...

OpenStudy (agent0smith):

lol you're welcome!

OpenStudy (lilai3):

hope you enjoy your medall!!